# Resource-Theoretic Hierarchy of Contextuality for General Probabilistic Theories

### v1

@@ 3.2 Operational theory associated to a GPT system
A preparation $P$ of a GPT system is a finite ensemble
$\bigl{(}\lambda_{i},\omega_{i}\bigr{)}_{i},$ (17)
where $(\lambda_{i})_{i}$ is a probability distribution. It has the same statistical behaviour as the GPT state
$\omega_{P}=\sum_{i}\lambda_{i}\omega_{i},$ (18)
which is called the mean of the preparation. Two preparations, $P=(\lambda_{i},\omega_{i})_{i}$ and $P^{\prime}=(\lambda^{\prime}_{i},\omega^{\prime}_{i})_{i}$, are said to be equivalent, denoted $P\sim P^{\prime}$, if their means are equal as GPT states. We denote the set of all preparations of a GPT system $\mathsf{A}$ by $\mathrm{Prep}(\mathsf{A})$.
Similarly, proofs of measurement contextuality require the notion of distinct measurement outcomes which are operationally equivalent. Let us now give the definition of measurements and measurement outcomes for GPT systems.
Definition 7.
A measurement $M$ of a GPT system $\mathsf{A}$ is a finite collection $\{e_{1},\ldots,e_{n}\}$ of effects that is normalized, i.e. it satisfies
$\sum_{k=1}^{n}e_{k}=1.$ (19)
The set of all finite ensembles of measurements of $\mathsf{A}$ is denoted by $\mathrm{Meas}(\mathsf{A})$, which has elements $(\lambda_{i},M_{i})_{i}$ where $(\lambda_{i})_{i}$ is a probability distribution.
Given such a measurement, the probability of observing outcome $k$ provided that system $\mathsf{A}$ is prepared in state $\omega$ is given by
$\mathrm{Pr}(k|M,\omega)=e_{k}\cdot\omega.$ (20)
Definition 8.
A measurement outcome $[k|M]$ of a GPT system is a pair of an effect $e_{k}$ and a measurement $M=\{e_{1},\ldots,e_{n}\}$ which contains $e_{k}$. To stress that the effect $e_{k}$ is thought of as part of the measurement $M$, we denote it by $e_{k|M}$. Two measurement outcomes $[k|M]$ and $[k^{\prime}|M^{\prime}]$ are equivalent, denoted $[k|M]\sim[k^{\prime}|M^{\prime}]$, whenever they are equal as GPT effects:
$e_{k|M}=e_{k^{\prime}|M^{\prime}}.$ (21)
The set of all ensembles of measurement outcomes of a GPT system $\mathsf{A}$ has elements $(\lambda_{i},e_{i})_{i}$, with $(\lambda)_{i}$ a probability distribution, and is denoted by $\mathrm{Out}(\mathsf{A})$. [^4] An ensemble of outcomes $\bigl{\{}\bigl{(}\lambda,[k|M]\bigr{)},\bigl{(}1-\lambda,[k^{\prime}|M^{\prime}]\bigr{)}\bigr{\}}$ is constructed by considering the ensemble of measurements $\{(\lambda,M),(1-\lambda,M^{\prime})\}$ and identifying the outcomes $[k|M]$ and $[k^{\prime}|M^{\prime}]$. The effect corresponding to this new outcome is $\lambda e_{k|M}+(1-\lambda)e_{k^{\prime}|M^{\prime}}$. The operational equivalences of the set can be defined in analogy to Definition 6.
Remark 9 (GPT system defined by an operational theory).
State and effect spaces defining a GPT system can be constructed from an operational theory specified by a collection of preparation and measurement procedures [59], which are list of instructions to be carried out in the lab, as well as a set of all possible conditional probabilities of a measurement outcome given a preparation. In this approach, states are defined to be equivalence classes of preparation procedures, where two preparation procedures are equivalent if they cannot be distinguished using any available measurement procedure. Similarly effects are defined as equivalence classes of measurement outcomes. While this construction is more general (for instance it might include preparation procedures with extraneous information), the only preparations considered in proofs of preparation contextuality are those that correspond to preparations as introduced in Definition 6 and similarly for proofs of measurement contextuality and measurement outcomes as defined in Definition 8.
One is often particularly interested in preparations supported on the extremal points of the convex set of states $\Omega$. For classical systems, extremal states are point distributions. For quantum systems, they are the rank 1 density operators. While in the classical case there are no non-trivial equivalences among preparations of extremal states, in quantum theory they do arise.
3.3 Simulations of GPT systems
Recall from Section 2 that a physical system is noncontextual if it admits of a noncontextual ontological model. An ontological model is traditionally defined for an operational theory, from which the GPT system can be derived, as we discussed in Remark 9. In this article, we associate a canonical operational theory to every GPT system, by endowing it with the space $\mathrm{Prep}(\mathsf{A})$ of preparation procedures, $\mathrm{Meas}(\mathsf{A})$ of measurement procedures and $\mathrm{Out}(\mathsf{A})$ of measurement outcomes.
This allows us to define an ontological model for a GPT system.
Definition 10.
Given a GPT system $\mathsf{A}=(\Omega_{\mathsf{A}},E_{\mathsf{A}},V_{\mathsf{A}})$ a discrete ontological model of $\mathsf{A}$ is given by a finite classical system $\Delta_{n}$ and maps
$\begin{split}\mu:\mathrm{Prep}(\mathsf{A})&\to\Delta_{n},\\ \mu:P&\mapsto\mu_{\mathcal{P}},\end{split}\qquad\qquad\begin{split}\chi:\mathrm{Out}(\mathsf{A})&\to\Delta_{n}^{*},\\ \chi:[k|M]&\mapsto\chi_{k|M_{j}}\end{split}$ (22)
satisfying
$\chi_{k|M}\cdot\mu_{P}=e_{k|M}\cdot\omega_{P}$ (23)
for any preparation $P\in\mathrm{Prep}(\mathsf{A})$ and any measurement outcome $[k|M]\in\mathrm{Out}(\mathsf{A})$, where $\omega_{P}$ is the mean of $P$. The ontological model is called preparation noncontextual if for any pair of equivalent preparations $P$ and $P^{\prime}$, we have
$\mu_{P}=\mu_{P^{\prime}}.$ (24)
Similarly the ontological model is called measurement noncontextual if for any pair of equaivalent outcomes $[k|M]$ and $[k^{\prime}|M^{\prime}]$ we have
$\chi_{k|M}=\chi_{k^{\prime}|M^{\prime}}.$ (25)
A model which is both measurement and preparation noncontextual is a noncontextual model.
Definition 11.
If $\mathsf{A}$ admits of a discrete noncontextual ontological model, we say that it is a noncontextual GPT system. We denote the collection of all such systems by $\mathfrak{G}_{\mathtt{NC}}$. A system for which all ontological models are contextual is called a contextual system.
A noncontextual ontological model represents the behaviour of the chosen system by the behaviour of a classical system, while preserving operational equivalences (Definitions 6 and 8). It is, in fact, a special case of the following notion of simulation [32].
Definition 12.
  • a multi-valued function $\Gamma:\Omega_{\mathsf{A}}\to\Omega_{\mathsf{B}}$ called state simulation and
  • a multi-valued function $\Theta:E_{\mathsf{A}}\to E_{\mathsf{B}}$ called effect simulation,
which is empirically adequate up to an error $\epsilon\in[0,\infty)$:
$\big{\lvert}e\cdot\omega-f\cdot\gamma\big{\rvert}\leq\epsilon\qquad\forall\,f\in\Theta(e),\;\forall\,\gamma\in\Gamma(\omega),$ (26)
and which preserves convex mixtures: [^5] Analogously, we require that the same holds for mixtures of effects rather than states, just as in [32].
$\lambda\Gamma(\omega)+(1-\lambda)\Gamma(\omega^{\prime})\subseteq\Gamma\bigl{(}\lambda\omega+(1-\lambda)\omega^{\prime}\bigr{)}\qquad\lambda\in[0,1]\text{ and }\omega,\omega^{\prime}\in\Omega_{\mathsf{A}},$ (27)
and the 00 effect:
$0_{\mathsf{B}}\in\Theta(0_{\mathsf{A}}).$ (28)
A 00-simulation is called an exact simulation.
Proposition 13 (Theorem 1 from [32]).
@@ 3.3 Simulations of GPT systems
A discrete ontological model of a GPT system $\mathsf{A}$ provides an exact simulation of $\mathsf{A}$ by a finite classical GPT system (and vice versa).
@@
Given a preparation $P=(\lambda_{i},\omega_{i})_{i}$ of the GPT system $\mathsf{A}$ and a simulation as above, preparations of $\mathsf{B}$ consistent with $P$ are the elements of the following subset of $\Omega_{\mathsf{B}}$:
[…]
Given an outcome $[k|M]=(\lambda_{i},e_{i})_{i}$ of the GPT system $\mathsf{A}$ and a simulation as above, outcomes of $\mathsf{B}$ consistent with $[k|M]$ are the elements of the following subset of $E_{\mathsf{B}}$:
[…]
We can use this notion to generalize the noncontextuality property of ontological models to the case of simulations following [32].
Definition 14.
A simulation of GPT systems is termed preparation univalent if for any two equivalent preparations $P$ and $P^{\prime}$, we have
$Q\sim Q^{\prime}$ (31)
for all $Q\in\Gamma(P)$ and all $Q^{\prime}\in\Gamma(P^{\prime})$.
A simulation of GPT systems is termed measurement univalent if for any two equivalent measurement outcomes $[k|M]$ and $k^{\prime}|M^{\prime}$, we have
$[q|W]\sim[q^{\prime}|W^{\prime}]$ (32)
for all $[q|W]\in\Theta([k|M])$ and all $[q^{\prime}|W^{\prime}]\in\Theta([k^{\prime}|M^{\prime}])$.
A simulation which is both preparation and measurement univalent is a univalent simulation.
@@
Given a bit $\Delta_{2}$ and a trit $\Delta_{3}$ one can define a univalent simulation $(\Gamma,\Theta)$ of $\Delta_{2}$ by $\Delta_{3}$ via a direct embedding. Namely, the state and effect simulations act on the bases of $\mathbb{R}^{2}$ and its dual (see Example 3) via
@@
and can be extended to all states and effects by linearity. This simulation is univalent since both $\Gamma$ and $\Theta$ are single-valued, a connection shown below in Proposition 18.
@@
[…]
@@
In [32] Müller and Garner define a univalent simulation to be single-valued. Our opinion is that Definition 14, stipulating preservation of operational equivalence, is more directly connected to generalized contextuality. Nevertheless, the two definitions are equivalent by Proposition 18 below. As a result, univalent $\epsilon$-simulations can be identified with the notion of $\epsilon$-embeddings [32, Definition 2], for which both the state and effect simulations are (single-valued) linear maps. Indeed, one can show that every univalent $\epsilon$-simulation extends to an essentially unique $\epsilon$-embedding and vice versa [32, Lemma 2].
Definition 17.
Given a pair of GPT systems $(\mathsf{A},\mathsf{B})$, we say that $\mathsf{A}$ is embeddable within $\mathsf{B}$, denoted by $\mathsf{A}\hookrightarrow\mathsf{B}$, if there exists an exact univalent simulation (equivalently a 00-embedding) of type $\mathsf{A}\to\mathsf{B}$. The relation $\hookrightarrow$ is called the embeddability preorder.
Indeed, one can easily show that $\hookrightarrow$ is a preorder relation. Specifically, the identity maps provide embeddings of type $\mathsf{A}\hookrightarrow\mathsf{A}$, which proves its reflexivity. To prove that it is a transitive relation, one can use the triangle inequality (see Proposition 21).
Proposition 18 (Theorem 1 [32]).
A simulation is preparation univalent if and only if its state simulation is single-valued. Similarly, a simulation is measurement univalent if and only if its effect simulation is single valued.
Proof.
First, assume that $\Gamma$ is single-valued. Then, for every preparation $P$, the set $\Gamma(P)$ has exactly one element, which is
$\bigl{(}\lambda_{i},\Gamma(\omega_{i})\bigr{)}.$ (36)
The simulation thus preserves preparation means by Equation 27, and is therefore preparation univalent.
On the other hand, consider any simulation that is not single-valued. That is, there is a state $\omega\in\Omega_{\mathsf{A}}$ such that there are two distinct states $\sigma_{1},\sigma_{2}\in\Omega_{\mathsf{B}}$ satisfying
$\{\sigma_{1},\sigma_{2}\}\subseteq\Gamma(\omega).$ (37)
Applying the state simulation to the preparation given by the delta distribution on $\omega$ gives
$\{\delta_{\sigma_{1}},\delta_{\sigma_{2}}\}\subseteq\Gamma(\delta_{\omega}).$ (38)
However, the means of the two preparations on the left are distinct — they are given by $\sigma_{1}$ and $\sigma_{2}$ respectively. Thus, the simulation is not preparation univalent.
An analogous proof holds for measurement univalence and single-valuedness of the effect simulation. ∎
@@ 3.4 Properties of univalent simulations
The following proposition, which stresses how the composition with a classical system does not affect the univalency of a simulation, is relevant for motivating our notion of contextuality preorder further on (Definition 23) as well as for quantifying contextuality.
@@
[…]
@@
Note that the value of excess is always in the interval $[0,1]$. One can show that it cannot exceed $1$ by observing that there is always a univalent $1$-simulation of type $\mathsf{A}\to\Delta_{1}$ and a univalent 00-simulation of type $\Delta_{1}\to\mathsf{B}$, and using the following proposition.
@@
$\big{\lvert}e_{\mathsf{A}}\cdot\omega_{\mathsf{A}}$ $-\beta\,\Theta(e_{\mathsf{A}})\cdot\alpha\,\Gamma(\omega_{\mathsf{A}})\big{\rvert}\leq$ (43)
@@
for any GPT system $\mathsf{A}$. A similar statement follows by setting the first term on the right-hand side of inequality (42) to zero. These two statements can be summarized as the following corollary of Proposition 21.
@@ 4 Hierarchy of contextuality
Recall that our interest in a hierarchy of contextuality of GPT systems stems from the wish to measure how useful contextuality is as a resource in information-theoretic applications. How can one decide whether a given measure of interest is indeed a well-motivated measure of contextuality? The resource-theoretic perspective [61, 26], fruitfully applied to studies of entanglement and many other resources in the past, suggests to first motivate a preorder $\succeq$ among the objects of interest. Importantly, some objects may be incomparable according to $\succeq$. In our case, the resource objects are GPT systems and the resource ordering is the advertised hierarchy of contextuality. The interpretation of such an ordering is that the relation $\mathsf{A}\succeq\mathsf{B}$ expresses the statement ‘$\mathsf{A}$ is at least as contextual as $\mathsf{B}$ is’. Given a good motivation for such a hierarchy, one can answer the above question as follows:
@@ 4.1 Motivation and the resource theory
    $\mathsf{A}\preceq\mathsf{A}\otimes\Delta_{n}\quad\text{ and }\quad\mathsf{A}\otimes\Delta_{n}\preceq\mathsf{A}$ (48)
    for every GPT system $\mathsf{A}$ and every classical GPT system $\Delta_{n}$. Here, $\otimes$ denotes the minimal composite from Definition 5.
@@
Notice that the first condition in Equation 48, i.e. $\mathsf{A}\preceq\mathsf{A}\otimes\Delta_{n}$, follows from desideratum 2 because every GPT system $\mathsf{A}$ is embeddable within $\mathsf{A}\otimes\Delta_{n}$ (see Proposition 19).
@@
    Consequently, using the fact that $\Delta_{1}\otimes\Delta_{n}$ is the same GPT system as $\Delta_{n}$, every classical GPT system is also at the bottom by Property (1).
@@
    Using definition 11 and desideratum 2, we thus get that all noncontextual GPT systems are equivalent to each other and lie at the bottom of the hierarchy.
Motivation for Requirement 1.
Contextuality, as we are thinking of it, is a property that expresses the degree to which a system’s behaviour escapes a purely classical explanation. In this sense, considering additional classical systems should not affect this property. An important point to note is that within the tensor product $\mathsf{A}\otimes\Delta_{n}$ (Definition 5), only correlations involving separable states between the generic GPT system $\mathsf{A}$ and the classical system are allowed. We can think of behaviours of $\mathsf{A}\otimes\Delta_{n}$ as dilations of those of $\mathsf{A}$ [62], given a classical environment variable $\Delta_{n}$.
If there were classical environmental variables, whose knowledge reduces the contextuality present in $\mathsf{A}$, then the contextuality of $\mathsf{A}$ would not be an authentic feature of $\mathsf{A}$, rather a consequence of one’s too narrow focus on $\mathsf{A}$ as opposed to its dilation $\mathsf{A}\otimes\Delta_{n}$. We can also support this argument on the technical side. Indeed, by Proposition 19, any noncontextual model of $\mathsf{A}\otimes\Delta_{n}$ provides also a noncontextual model of $\mathsf{A}$. In other words, $\mathsf{A}$ is noncontextual if and only if $\mathsf{A}\otimes\Delta_{n}$ is. [^6] Note that under the minimal tensor product of Definition 5 allowing access to arbitrary noncontextual systems (and not just classical systems) would lead to the same hierarchy, which follows from the fact that access to classical systems already places all noncontextual systems at the bottom of the hierarchy. The hierarchy of contextuality ought to capture this fact.
Motivation for Requirement 2.
We can also think of contextuality (of a GPT system) as the inability to provide a noncontextual model (Definition 11) for its statistical behaviours. If a GPT system $\mathsf{A}$ can be exactly simulated by a GPT system $\mathsf{B}$ via a univalent simulation, then any obstruction to constructing such a noncontextual model for $\mathsf{A}$ must already be present in $\mathsf{B}$. In this sense, we think of $\mathsf{B}$ as being “at least as contextual” as $\mathsf{A}$ is. Indeed, provided with a noncontextual model of $\mathsf{B}$, i.e. an exact univalent simulation of $\mathsf{B}$ by a classical GPT system, one can construct a noncontextual model of $\mathsf{A}$ by composition with the simulation of $\mathsf{A}$ by $\mathsf{B}$.
@@
$\mathsf{A}\preceq\mathsf{B}\quad\iff\quad\exists\,n\in\mathbb{N}\text{ such that }\mathsf{A}\hookrightarrow\mathsf{B}\otimes\Delta_{n}\text{ holds.}$. (49)
@@
The resource theory of GPT-contextuality is defined as follows. Its objects are finite-dimensional GPT systems and each transformation is an exact simulation ${(\Gamma,\Theta):\mathsf{A}\to\mathsf{B}\otimes\Delta_{n}}$ for some $\mathsf{A},\mathsf{B}\in\mathfrak{G}_{\mathrm{fin}}$. It acts on objects by transforming $\mathsf{B}$ to $\mathsf{A}$ and is undefined for any GPT system other than $\mathsf{A}$. Transformations
@@
can be composed (sequentially) to produce a transformation
$(\Gamma\,\alpha,\Theta\,\beta):\mathsf{A}\to\mathsf{C}\otimes\Delta_{m\cdot k}$ (51)
by using the isomorphism between $\Delta_{m}\otimes\Delta_{k}$ and $\Delta_{m\cdot k}$. Free transformations are those ones that are also univalent simulations.
This gives rise to a concrete resource theory in the sense of [26, Definition 3.10]. Moreover, the resulting resource ordering coincides with our hierarchy of contextuality. Note that in a resource theory the existence of a free operation transforming a resource object $r$ to a resource object $s$ defines the resource ordering denoted by $s\preceq r$. Thus, to match our definition of the hierarchy of contextuality, the transformation as in (50) has to be understood as a free operation from $\mathsf{B}$ to $\mathsf{A}$ (and not from $\mathsf{A}$ to $\mathsf{B}$ as the notation might suggest). This is consistent with the interpretation of a simulation of $\mathsf{A}$ by $\mathsf{B}$ whereby an experimenter has access to $\mathsf{B}$ and uses it to simulate $\mathsf{A}$, hence transforming their access to system $\mathsf{B}$ to their effective access to system $\mathsf{A}$.

### v2

@@ 3.3 Simulations of GPT systems
Given a pair of GPT systems $(\mathsf{A},\mathsf{B})$, we say that $\mathsf{A}$ is embeddable within $\mathsf{B}$, denoted by $\mathsf{A}\hookrightarrow\mathsf{B}$, if there exists an exact univalent simulation of type $\mathsf{A}\to\mathsf{B}$. The relation $\hookrightarrow$ is called the embeddability preorder.
Indeed, one can easily show that $\hookrightarrow$ is a preorder relation. Specifically, identity maps provide univalent simulations of type $\mathsf{A}\hookrightarrow\mathsf{A}$, which proves the reflexivity of $\hookrightarrow$. To prove that it is a transitive relation, one can use the triangle inequality (see Proposition˜12).
It has also been shown that
@@
  • a discrete ontological model of a GPT system $\mathsf{A}$ provides an exact simulation of $\mathsf{A}$ by a finite classical GPT system (and vice versa) [64, Theorem 1],
  • a discrete noncontextual ontological model of a GPT system $\mathsf{A}$ is the same as an exact univalent simulation of $\mathsf{A}$ by a finite classical GPT system [64, Corollary 1].
In this sense (univalent) simulations generalize (noncontextual) ontological models.
@@
Given a bit $\Delta_{2}$ and a trit $\Delta_{3}$ one can define a univalent simulation $(\Gamma,\Theta)$ of $\Delta_{2}$ by $\Delta_{3}$ via a direct embedding. Namely, the state and effect simulations act on the bases of $\mathbb{R}^{2}$ and its dual (see Example˜3) via
@@
and can be extended to all states and effects by linearity. This simulation is univalent since both $\Gamma$ and $\Theta$ are single-valued.
@@
$\Gamma\left(\delta_{2}\right)=\big\{\lambda\delta_{2}+(1-\lambda)\delta_{3}\nonscript\;\,\big|\allowbreak\nonscript\;\,\mathopen{}\lambda\in[0,1]\big\},\qquad\Theta\left(\delta_{2}^{*}\right)=\left\{\delta_{2}^{*}+\delta_{3}^{*}\right\}$ (22)
@@ 3.4 Properties of univalent simulations
The following proposition, which stresses how the composition with a classical system does not affect the univalency of a simulation, is relevant for motivating our notion of contextuality preorder further on (Definition˜14) as well as for quantifying contextuality.
@@
$\varepsilon(A,B)\coloneqq\inf\big\{\epsilon\in[0,\infty)\nonscript\;\,\big|\allowbreak\nonscript\;\,\mathopen{}\text{there is a univalent $\epsilon$-simulation $A\to B$}\big\},$ (26)
@@
Note that the value of excess is always in the interval $[0,1]$. One can show that it cannot exceed $1$ by observing that there is always a univalent $1$-simulation of type $\mathsf{A}\to\Delta_{1}$ and a univalent 0-simulation of type $\Delta_{1}\to\mathsf{B}$, and using the following proposition.
@@
$\big\lvert e_{\mathsf{A}}\cdot\omega_{\mathsf{A}}$ $-\beta\,\Theta(e_{\mathsf{A}})\cdot\alpha\,\Gamma(\omega_{\mathsf{A}})\big\rvert$ (28)
@@
for any GPT system $\mathsf{A}$. A similar statement follows by setting the first term on the right-hand side of inequality (27) to zero. These two statements can be summarized as the following corollary of Proposition˜12.
@@ 4 Hierarchy of contextuality
Recall that our interest in a hierarchy of contextuality of GPT systems stems from the wish to measure how useful contextuality is as a resource in information-theoretic applications. How can one decide whether a given measure of interest is indeed a well-motivated measure of contextuality? The resource-theoretic perspective [25, 36], fruitfully applied to studies of entanglement and many other resources in the past, suggests to first motivate a preorder $\succeq$ among the objects of interest. Importantly, some objects may be incomparable according to $\succeq$. In our case, the resource objects are GPT systems and the resource ordering is the advertised hierarchy of contextuality. The interpretation of such an ordering is that the relation $\mathsf{A}\succeq\mathsf{B}$ expresses the statement $\mathsf{A}$ is at least as contextual as $\mathsf{B}$ is’. Given a good motivation for such a hierarchy, one can answer the above question as follows:
@@ 4.1 Motivation and the resource theory
[…]
    for every GPT system $\mathsf{A}$ and every classical GPT system $\Delta_{n}$. Here, $\otimes$ denotes the minimal composite from Definition˜5.
@@
Notice that the first condition in Equation˜33, i.e. $\mathsf{A}\preceq\mathsf{A}\otimes\Delta_{n}$, follows from desideratum 2 because every GPT system $\mathsf{A}$ is embeddable within $\mathsf{A}\otimes\Delta_{n}$ (see Proposition˜10).
@@
    Consequently, using the fact that $\Delta_{1}\otimes\Delta_{n}$ is the same GPT system as $\Delta_{n}$, every classical GPT system is also at the bottom by Desideratum 1.
@@
    Finally, using the fact that noncontextual systems are simplex embeddable and Desideratum 2, we conclude that all noncontextual GPT systems are equivalent to each other and lie at the bottom of the hierarchy.
Motivation for Desideratum 1.
Contextuality, as we are thinking of it, is a property that expresses the degree to which a system’s behaviour escapes a purely classical explanation. In this sense, considering additional classical systems should not affect this property. An important point to note is that within the tensor product $\mathsf{A}\otimes\Delta_{n}$ (Definition˜5), only correlations involving separable states between the generic GPT system $\mathsf{A}$ and the classical system are allowed. We can think of behaviours of $\mathsf{A}\otimes\Delta_{n}$ as dilations of those of $\mathsf{A}$ [47], given a classical environment variable $\Delta_{n}$.
If there were classical environmental variables, whose knowledge reduces the contextuality present in $\mathsf{A}$, then the contextuality of $\mathsf{A}$ would not be an authentic feature of $\mathsf{A}$, rather a consequence of one’s too narrow focus on $\mathsf{A}$ as opposed to its dilation $\mathsf{A}\otimes\Delta_{n}$. We can also support this argument on the technical side. Indeed, by Proposition˜10, any noncontextual model of $\mathsf{A}\otimes\Delta_{n}$ provides also a noncontextual model of $\mathsf{A}$. In other words, $\mathsf{A}$ is noncontextual if and only if $\mathsf{A}\otimes\Delta_{n}$ is. [^7] Note that under the minimal tensor product of Definition 5 allowing access to arbitrary noncontextual systems (and not just classical systems) would lead to the same hierarchy, which follows from the fact that access to classical systems already places all noncontextual systems at the bottom of the hierarchy. The hierarchy of contextuality ought to capture this fact.
Motivation for Desideratum 2.
We can also think of contextuality (of a GPT system) as the inability to provide a noncontextual model (i.e. a univalent simulation by a classical system) for its statistical behaviours. If a GPT system $\mathsf{A}$ can be exactly simulated by a GPT system $\mathsf{B}$ via a univalent simulation, then any obstruction to constructing such a noncontextual model for $\mathsf{A}$ must already be present in $\mathsf{B}$. In this sense, we think of $\mathsf{B}$ as being “at least as contextual” as $\mathsf{A}$ is. Indeed, provided with a noncontextual model of $\mathsf{B}$, i.e. an exact univalent simulation of $\mathsf{B}$ by a classical GPT system, one can construct a noncontextual model of $\mathsf{A}$ by composition with the simulation of $\mathsf{A}$ by $\mathsf{B}$.
@@
$\mathsf{A}\preceq\mathsf{B}\quad\iff\quad\exists\,n\in\mathbb{N}\text{ such that }\mathsf{A}\hookrightarrow\mathsf{B}\otimes\Delta_{n}\text{ holds.}$ (34)
@@
The resource theory of GPT-contextuality is defined as follows. Its objects are finite-dimensional GPT systems in $\mathfrak{G}_{\mathrm{fin}}$. The set of transformation from $\mathsf{B}$ to $\mathsf{A}$ is identified with exact simulations of type $\mathsf{A}\to\mathsf{B}\otimes\Delta_{n}$}. Two such simulations
@@
can be composed (sequentially) to produce a simulation
$(\alpha\,\Gamma,\beta\,\Theta):\mathsf{A}\to\mathsf{C}\otimes\Delta_{m\cdot k}$ (36)
via the canonical isomorphism between $\Delta_{m}\otimes\Delta_{k}$ and $\Delta_{m\cdot k}$. Free transformations are those simulations that are also univalent.
This gives rise to a concrete resource theory in the sense of [36, Definition 3.10]. Moreover, the resulting resource ordering coincides with our hierarchy of contextuality. In a general resource theory, the existence of a free transformation from a resource object $r$ to a resource object $s$ defines the resource ordering denoted by $s\preceq r$, i.e. $r$ is at least as good as a resource than $s$ is.
In our work, a simulation of type $\mathsf{C}\to\mathsf{D}$ is interpreted as providing effective access to system $\mathsf{C}$ to an agent with access to $\mathsf{D}$. In this sense, access to $\mathsf{D}$ grants the agent at least as many capabilities as access to $\mathsf{C}$. For this reason, a simulation of type $\mathsf{A}\to\mathsf{B}\otimes\Delta_{n}$ is identified with a resource-theoretic transformation from $\mathsf{B}$ to $\mathsf{A}$ and not vice versa. Note, however, that it is not in general realisable by a physical transformation from $\mathsf{B}$ to $\mathsf{A}$ in the GPT sense. A physical transformation $\mathsf{B}\to\mathsf{A}$ is given by a linear map $M:V_{\mathsf{B}}\to V_{\mathsf{A}}$ such that $M(\Omega_{\mathsf{B}})\subseteq\Omega_{\mathsf{A}}$ and $M^{*}(E_{\mathsf{A}})\subseteq E_{\mathsf{B}}$. For further discussion see Sections˜5.1 and A.